OpenStudy (anonymous):

What are the two arithmetic means between -13 and 8?

OpenStudy (mayankdevnani):

first arithmetic mean is average. \[\huge Average=\frac{-13+8}{2}=\frac{-5}{2}=-2.5\]

OpenStudy (anonymous):

My options are: -5, 2 -6, 1 -4, 3 -3, 4

OpenStudy (mayankdevnani):

i think its option A

OpenStudy (anonymous):

How do you arrive at it?

OpenStudy (mayankdevnani):

because see my fraction involves both -5 and 2. that's why i choose option a

OpenStudy (hunus):

I don't think they are actually talking about arithmetic means or there wouldn't be two. I think they're talking about a sequence where there are two terms between -13 and 8 \[\Large -13,a_{2},a_{3},8\] If you use the formula for an arithmetic progression \[\Large a_{n}=a_{1}+(n-1)d\]Where d is the common difference and n is the term you're on \[\Large a_{4}=8=-13+3d\] So that d is 7, then the two middle terms would be \[\Large a_{2}=a_{1}+d=-13+7=-6\]\[\Large a_{3}=a_{1}+2d=-13+14=1\]

OpenStudy (anonymous):

Alright, thanks!

OpenStudy (hunus):

Yup :) Understand the steps?